# The syzygy distinguisher

### v2

@@ Abstract
We present a new distinguisher for alternant and Goppa codes, whose complexity is subexponential in the error-correcting capability. It does not suffer from the strong regime limitations of the previous distinguishers or structure recovery algorithms: in particular, it applies to the codes used in the Classic McEliece candidate for postquantum cryptography standardization. The invariants that allow us to distinguish are graded Betti numbers of the homogeneous coordinate ring of a shortening of the dual code.
@@ 1 Introduction
Assumption 2, on the other hand, stands on a firm theoretical ground: the decoding problem for generic linear codes is known to be NP-hard [2]. As such, it is believed to resist the advent of quantum computers, which made the McEliece system a good candidate for postquantum cryptography. However, starting with Prange’s information set decoding algorithm [24], generic decoding methods saw continuous incremental improvements over time. Joint with technological progress in computational power, this eventually led to practical message recovery attacks against the McEliece system with its initially proposed parameter set. However, it appears that this weakness was only the result of a too optimistic choice of parameters. With the need for new standards for postquantum cryptography, an updated version named Classic McEliece was proposed, still relying on binary Goppa codes, but with more conservative parameter sets. In the design rationale of this new system [4] one can find an impressive list of several dozens of papers on generic decoding algorithms, ranging over the last five decades. It is then observed that all these algorithms have complexity exponential in the error-correcting capability of the code. Better, the constant in this exponential is still the same as in Prange’s original result: all improvements remain confined in terms of lesser order! This gives a quite convincing feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
This is false: in this work, we present a new distinguisher for alternant and Goppa codes, whose complexity is subexponential in the error-correcting capability. It does not suffer from the strong regime limitations of the previous distinguishers or key recovery methods: in particular, it applies to the codes used in Classic McEliece.
@@
Syzygies sometimes appear as a tool in cryptanalytic works, or in the study of Gröbner basis algorithms; however in general these are only syzygies of low order. Likewise, a few works in coding theory (such as [26] or [15]) use homological properties of finite sets of points; but the applications differ from ours.
@@ Acknowledgements
The author learned a lot from the code-based cryptography seminar organized by Jean-Pierre Tillich at Inria Paris. More important perhaps, without his friendly pressure, probably this work would have stayed eternally in limbo. He should be thanked for all that. The very favourable environment towards research at ANSSI in general, and at LCR in particular, was also highly helpful. Henri Gilbert’s advices led to improvements in several aspects of the presentation.
@@ 2 Minimal resolutions and graded Betti numbers
Then the the minimal resolution of ${\mathsf{C}}^{\langle\cdot\rangle}$ canonically admits the direct sum of the linear strands of the minimal resolutions of $S/\langle V^{(1)}\rangle$, …, $S/\langle V^{(l)}\rangle$ as a subcomplex. In particular for all $r\geq 2$ we have
@@ 4 Regularity $2$ and the small defect heuristic
  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $13\leq r\leq k+$ 1.
@@ 5 The distinguisher, with and without shortening
$\kappa=\widetilde{O}\left(q^{\omega\frac{R^{2}}{1-R}\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}\right).$ (93)
Observe [5, §3.4] that the security parameter in the Classic McEliece system is linear in $t\propto n/\log(n)$. Then in (93) we have $\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$, which means that this complexity is subexponential in the security parameter.
@@ 6 Conclusion and open problems
We presented the first analysis of the McEliece cryptosystem whose complexity is subexponential in the security parameter. However this is only an asymptotic result. For concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the security parameter by a non-negligible factor.
@@
Pursue this study in a systematic way. This should cover basic operations on codes that have a standard geometric interpretation, such as puncturing (discarding one subset from a set of points), shortening (projection from one or several of the points), duality of codes (“Gale duality”), tensor product of codes (Segre embedding), and powers of codes (Veronese embedding), but also other operations such as the $(u|u+v)$ construction, and crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.

### v3

@@ Abstract
We present a new distinguisher for alternant and Goppa codes, whose complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover it does not suffer from the strong regime limitations of the previous distinguishers or structure recovery algorithms: in particular, it applies to the codes used in the Classic McEliece candidate for postquantum cryptography standardization. The invariants that allow us to distinguish are graded Betti numbers of the homogeneous coordinate ring of a shortening of the dual code.
@@ 1 Introduction
Assumption 2, on the other hand, stands on a firm theoretical ground: the decoding problem for generic linear codes is known to be NP-hard [2]. As such, it is believed to resist the advent of quantum computers, which made the McEliece system a good candidate for postquantum cryptography. However, starting with Prange’s information set decoding algorithm [24], generic decoding methods saw continuous incremental improvements over time. Joint with technological progress in computational power, this eventually led to practical message recovery attacks against the McEliece system with its initially proposed parameter set. However, it appears that this weakness was only the result of a too optimistic choice of parameters. With the need for new standards for postquantum cryptography, an updated version named Classic McEliece was proposed, still relying on binary Goppa codes, but with more conservative parameter sets adapted to resist the best generic decoding attacks with some margin of safety. In the design rationale of this new system [4] one can find an impressive list of several dozens of papers on generic decoding algorithms, ranging over the last five decades. It is then observed that all these algorithms have complexity exponential in the error-correcting capability of the code. Better, the constant in this exponential is still the same as in Prange’s original result: all improvements remain confined in terms of lesser order! This gives a quite convincing feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
This belief is false, and there is nothing special with Goppa codes. In this work, we present a new distinguisher for alternant and Goppa codes, i.e. a basic structural analysis of the McEliece cryptosystem, whose complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover it does not suffer from the strong regime limitations of the previous distinguishers or key recovery methods: in particular, it applies to the codes used in Classic McEliece.
@@
Syzygies sometimes appear as a tool in cryptanalytic works, or in the study of Gröbner basis algorithms; however in general only the first module of syzygies is considered, not those of higher order. Likewise, a few works in coding theory (such as [26] or [15]) use homological properties of finite sets of points; but the applications differ from ours.
@@ Acknowledgements
The author learned a lot from the code-based cryptography seminar organized by Jean-Pierre Tillich at Inria Paris. More important perhaps, without his friendly pressure, probably this work would have stayed eternally in limbo. He should be thanked for all that. The very favourable environment towards research at ANSSI in general, and at LCR in particular, was also highly helpful. Henri Gilbert’s advices led to improvements in several aspects of the presentation. Anonymous members of LCR unsuccessfully suggested the title “The Dyztynguysher” for this work.
@@ 2 Minimal resolutions and graded Betti numbers
Then the minimal resolution of ${\mathsf{C}}^{\langle\cdot\rangle}$ canonically admits the direct sum of the linear strands of the minimal resolutions of $S/\langle V^{(1)}\rangle$, …, $S/\langle V^{(l)}\rangle$ as a subcomplex. In particular for all $r\geq 2$ we have
@@ 4 Regularity $2$ and the small defect heuristic
  • $\beta_{r-2,r}=(r-2)\binom{k-1}{r-2}\;$ for $3\leq r\leq k+1\quad$ (so $\beta_{r-2,r}=\beta_{k+2-r,k+4-r}$).
@@ 5 The distinguisher, with and without shortening
$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}.$ (93)
Observe [5, §3.4] that the security parameter in the Classic McEliece system, based on the complexity of generic decoding algorithms, is linear in $t\propto n/\log(n)$. Then in (93) we have $\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$, which means that our complexity is subexponential in the security parameter.
@@ 6 Conclusion and open problems
We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms, more precisely, subexponential in the error-correcting capability of the code. However this is only an asymptotic result. For concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the security parameter by a non-negligible factor.
@@
Pursue this study in a systematic way. This should cover basic operations on codes that have a standard geometric interpretation, such as puncturing (discarding one subset from a set of points), shortening (projection from one or several of the points), duality of codes (“Gale duality”), tensor product of codes (Segre embedding), and powers of codes (Veronese embedding), but also other operations such as the $(u|u+v)$, subfield subcode, or trace code constructions; and crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.